Bordon is ten years older than his sister. In three years, he will be twice as old as she is now. How old are they now?
Answer by mananth(16949)

Bordon is ten years older than his sister. In three years, he will be twice as old as she is now. How old are they now?
let sister's age be x
Bordon's age will be x+10
After three years

And Bordens age will be x+10+3
x+10+3= 2x
10+3= 2x-x
x=13------------- sister's age
Bordon's age ---------- 13+10= 23 years